What is the mass, in milligrams, of one gold atom? Express your answer in scientific notation to 3 significant figures
1
Expert's answer
2018-03-31T08:54:08-0400
m(Au)=1.66*10⁻²⁴⁺³*196.9665=3.27*10⁻²²